# An Ultra-Sub-Wavelength Microwave Polarization Switch Implemented with Directed Surface Acoustic Waves in a Magnonic Crystal

### v2

@@ Abstract
The ability to switch the polarization of a transmitted electromagnetic wave from vertical to horizontal, or vice versa, is of great technological interest because of its many applications in long distance communication. Binary bits can be encoded in two orthogonal polarizations and transmitted securely from point to point. Polarization switches, however, are usually much larger than the wavelength of the electromagnetic wave. Consequently, most research in this area has focused on the optical regime where the wavelength is relatively short (1 m), so that the switch being much larger than the wavelength is not too inconvenient. However, this changes in the microwave regime where the wavelength is much larger (typically 1 cm). That makes a microwave ultra-sub-wavelength polarization switch very attractive. Here, we report such a switch made of an array of magnetostrictive nanomagnets (100 nm lateral dimension) deposited on a piezoelectric substrate to make an “artificial magnonic crystal”. A surface acoustic wave (SAW) launched in the substrate with suitable electrodes excites spin waves in the nanomagnets via phonon-magnon coupling, resulting in radiation of electromagnetic waves via magnon-photon coupling. The polarization of the beam radiated in one particular direction at a given frequency can be rotated through 90∘ by switching the direction of SAW propagation in the piezoelectric substrate between two mutually orthogonal directions. By aligning the transmitter and the receiver along that particular direction (known only to authorized users), one can communicate securely from point to point, without the need for encryption or cryptography.
@@
polarization switch, ultra-sub-wavelength magneto-elastic antenna, magnonic crystal, surface acoustic wave1 Introduction
@@
For digital data transmission via two orthogonal polarizations encoding bits 0 and 1, one would require a polarization switch that will switch the polarization of an electromagnetic wave from approximately horizontal (encoding the bit 0) to approximately vertical (encoding the bit 1), or vice versa. Polarization purity is not a concern for such applications. As long as the two polarizations are distinguishable (i.e., approximately orthogonal), it will suffice. Most polarization switches work at optical frequencies and are implemented with various techniques such as rotating waveplates [3], Babinet-Soleil compensators [4], Berek rotary compensators [5], fiber coil polarization controllers [6], Faraday rotators [7], degree of polarization generators [8], lithium niobate electro-optics [9], liquid crystals [10], digital micromirrors [11], graphene metasurfaces [12] and on-chip photonic circuits [13, 14, 15].
All of the above constructions are typically much larger than the wavelength. While this is less of a problem in the domain of optics where the wavelength is usually 1 -10 m, it is a serious problem in the microwave frequency region where the wavelength can be several cm. Therefore, sub-wavelength polarization switches are very attractive for the microwave regime. Here, we demonstrate a microwave polarization switch (working at 1-30 GHz) that is orders of magnitude smaller than the wavelength. This can have applications in ultra-compact polarization division multiplexers and digital data transmitters in the microwave range.
@@ 2 Results and Discussion
The structure of the polarization switch is shown in Fig. 1. It consists of a periodic two-dimensional array of magnetostrictive nanomagnets (made of cobalt) deposited on a piezoelectric substrate (LiNbO3). This is a periodic system for a spin wave (magnons) in the nanomagnets and hence it is called an “artificial magnonic crystal”. It is the same system that was used in ref. [17] for beam steering. Here, it is used for a different purpose, namely polarization switching. All antenna properties such as the scattering parameter spectrum, radiation efficiencies at different frequencies, etc. can be found in refs. [16, 17].
@@
The LiNbO3 substrate is 0.5 mm thick. The nanomagnets are slightly elliptical with major axis 110 nm and minor axis 100 nm. The thickness is 6 nm. Each nanomagnet has a 5 nm thick Ti layer underneath for adhesion to the substrate. The nanomagnet array covers an area of 100 m 100 m and is fabricated with electron-beam-lithography using a Raith Voyager e-beam writer. The substrate is spin-coated with a single layer of PMMA resist spun at 2500 rpm and baked at 110∘C for 2 minutes. It is then patterned with e-beam, and developed in a methyl isobutyl ketone and isopropyl alcohol (MIBK-IPA, 1:3) solution for 60 seconds, followed by a cold IPA rinse. After the patterning is complete, a 5 nm-thick titanium (Ti) adhesion layer is deposited using electron beam evaporation at a base pressure of 2.3 10-7 Torr, followed by the deposition of 6 nm thick cobalt. Lift-off is performed using remover PG solution. A scanning electron micrograph of the nanomagnets can be found in ref. [16]. The electrodes for launching surface acoustic waves are delineated with optical lithography.
@@
A surface acoustic wave (SAW) is launched in two mutually orthogonal directions by activating electrode pairs (3,4) or (5,6) (see Fig. 1). The SAW excites spin waves in the magnonic crystal via phonon-magnon coupling, which radiate electromagnetic waves via magnon-photon coupling [16, 17, 18]. Unlike a conventional electromagnetic antenna that radiates electromagnetic waves owing to fluctuating charges or time-varying electric dipoles, these structures (nano-antennas) radiate electromagnetic waves due to fluctuating magnetization associated with spin waves [16, 17, 18, 19, 20, 21]. Hence, they do not behave as traditional antennas and are not constrained by the Harrington limit or Chu’s limit that afflict the gain and bandwidth of traditional antennas.
@@
In the past, we measured the radiation pattern of these samples by propagating a SAW in two different directions [16]. We can apply a microwave frequency voltage between either electrode pairs (3,4) or (5,6) to launch SAWs in two mutually perpendicular directions. We then measure the radiation patterns in an anechoic chamber for these two different directions of SAW propagation. The patterns are measured in the plane of the nanomagnets and in the two planes transverse to the plane of the nanomagnets. The in-plane results are shown in Fig. 2 for both horizontal and vertical polarizations of the emitted radiation. The radiation patterns were measured in an AMS-8701 Anechoic Chamber, Antenna Measurement System using a 3164-10 Open Boundary Quad-ridged Horn Antenna.
@@
Fig. 2 shows that there is significant anisotropy in the radiation pattern at all frequencies measured (despite this being an ultra-sub-wavelength structure that should behave like a point source), but more importantly, the radiation pattern depends on the direction of SAW propagation (the difference between “orientation 1” and “orientation 2”). The difference is, of course, more pronounced at certain frequencies.
@@
We investigated the directional dependence via micromagnetic simulations. Theoretical simulations revealed that the nature of the spin waves excited in the nanomagnets is strongly influenced by whether the surface acoustic wave propagates parallel to the major axis of the elliptical nanomagnets (as would be the case if electrodes 3 and 4 in Fig. 1 were excited) or parallel to the minor axis of the nanomagnets (as would be the case if electrodes 5 and 6 were excited instead; see supplementary material of [17]). As a result, the electromagnetic radiation pattern also depends strongly on the direction of SAW propagation [16].
@@
Why the spin wave pattern depends on the direction of SAW propagation is easily understood. The surface acoustic wave subjects the nanomagnets to periodic strain. Strain acts like an effective magnetic field in a magnetostrictive nanomagnet and in the case of SAW, an effective periodic magnetic field will appear in the nanomagnet [22]. This periodic field will make the magnetization process and oscillate in time, thereby producing a spin wave. The direction of the effective magnetic field is roughly along the direction of SAW propagation [22]. Therefore, if we change the direction of SAW propagation, we will change the direction of the effective time-varying magnetic field in the nanomagnet and this will change the axis of precession of the magnetization within the nanomagnet, resulting in a change in the oscillations (both amplitude and phase) of the x-, y- and z-components of the magnetization, i.e. the spin wave pattern will change [17]. Ref. [17] found that this changed the radiation spectrum, whereas ref.[16] showed that this also changes the radiation pattern for either horizontal or vertical polarization.
@@
Since the spin waves radiate the electromagnetic waves, changing the spin wave patterns in space by changing the direction of SAW propagation also changes the polarization of the radiated beam in any direction. This polarization change was not studied earlier and here we report that study. The Supporting Information provides a phenomenological theory of why the polarization will change if we chnage the direction of SAW propagation.
@@ 2.1 Polarization dependence on the direction of SAW propagation
In Fig. 3, we first plot the radiation patterns in the plane of the nanomagnets for both vertical and horizontal polarizations in absolute units (as opposed to in db) for both directions of SAW propagation. This is done for different frequencies.
Let us say that at a given frequency the horizontal polarization value in a given direction at a given frequency is and the vertical polarization value is . We define a polarization angle as
@@
In Fig. 4, we show the radial plot of at different frequencies in radians. Note that at 5.9 GHz, in the 75∘ direction, 1.5 radians for orientation 2 and 0.1 radians for orientation 1. Hence, the beam is 99.5% horizontally polarized for orientation 1 and 99.5% vertically polarized for orientation 2. Thus, by switching the direction of SAW propagation from orientation 1 to orientation 2, we can change the polarization of the beam radiated in the 75∘ direction at 5.9 GHz (in the nanomagnets’ plane) from nearly horizontal to nearly vertical with very high degree of polarization purity, more than enough purity for binary encoding of data. We can place the receiver at 75∘ angle as shown in the right panel of Fig. 4 and receive either horizontal polarization or vertical polarization by switching the direction of SAW propagation by switching the electrode pairs that are activated to launch the SAW. This is the basis of the polarization switch. Note that the positions of the transmitting device and the receiver do not have to be fixed, but their relative alignment have to be fixed in this scheme. This enables point-to-point communication via polarization.
2.2 Security of communication
@@
The fact that the point-to-point communication via polarization encoding works only for a specific alignment of the transmitter and receiver adds an additional layer of security. The authorized sender has access to the polarization radial plot and can inform the authorized receiver where to place the receiving device. An unauthorized receiver, intent on intercepting the message (eavesdropping), will not know where to place the receiver and will place it at a wrong location with very high probability. The eavesdropper will therefore not receive binary polarization states and hence cannot eavesdrop or intercept the message. For example, if the eavesdropper places the receiver at 0∘ angle, then for orientation 1, 0 radians which means that the beam is horizontally polarized and for orientation 2, is again 0 radians and the beam is again horizontally polarized. If horizontal polarization encodes the bit 0, then the eavesdropper will constantly received a string of 0-s and no real message. Similarly, if the eavesdropper places the receiver at 180∘ angle, = 0.7 radians for orientation 1 and 1.3 radians for orientation 2. In this case, the received beam will be 58% horizontally polarized and 42% vertically polarized for orientation 1, while being 7% horizontally polarized and 93% vertically polarized for orientation 2. The poor polarization purity precludes deciphering the binary bits and hence the message. Thus, without precise knowledge of the correct location for the receiver, eavesdropping is impossible and confidentiality is assured. This eliminates the need for encrypting the message or resorting to cryptography, which are always vulnerable to sophisticated attacks.
@@
The scheme is also unclonable since the specific alignment that works will vary from sample to sample because of unavoidable manufacturing variations. Thus, the specific alignment is a “fingerprint” of any chosen antenna and hence can act as a physically unclonable function (PUF) that cannot be reproduced or predicted, thereby providing strong security.

### v3

# An Ultra-Sub-Wavelength Microwave Polarization Switching Antenna for Covert Communication Implemented with Surface Acoustic Waves in an Artificial Multiferroic Magnonic Crystal

@@ Abstract
The ability to switch at will the polarization of a transmitted electromagnetic wave from vertical to horizontal, or vice versa, is of great technological interest because of its many applications in long distance communication (e.g., polarization division multiplexing). Binary bits can be encoded in the two orthogonal polarizations and transmitted from point to point. Polarization switches, however, are usually much larger than the wavelength of the electromagnetic wave that they transmit. Consequently, most research in this area has focused on the optical regime where the wavelength is relatively short (1 m), so that the switch being much larger than the wavelength is not too inconvenient. However, this changes in the microwave regime where the wavelength is much larger (typically 1 cm). That makes a microwave ultra-sub-wavelength polarization switch very attractive. Here, we report, for the first time to our knowledge, such a switch made of an array of magnetostrictive nanomagnets (100 nm lateral dimension, 5 nm thickness) deposited on a piezoelectric substrate to make an “artificial multiferroic magnonic crystal (AMMC)”. A surface acoustic wave (SAW) launched in the substrate with suitable electrodes excites confined spin waves in the nanomagnets via phonon-magnon coupling, which then radiate electromagnetic waves in space via magnon-photon coupling. In some particular direction(s), determined by the AMMC parameters, the polarization of the beam at a given frequency can be rotated through 90∘ by switching the direction of SAW propagation in the piezoelectric substrate between two mutually orthogonal directions via activation of two different pairs of launching electrodes. By aligning the transmitter and the receiver along such a direction (known only to authorized users), one can communicate covertly from point to point, without the need for encryption or cryptography. Furthermore, this attribute also makes the antenna “stealthy” since the message can be concealed from any eavesdropper whose receiver is not precisely aligned in the correct direction.
@@
polarization switch, ultra-sub-wavelength magneto-elastic antenna, covert communication, stealthy antenna, multiple-input-multiple-output (MIMO) antenna, physical unclonable function1 Introduction
@@
For digital data transmission via two orthogonal polarizations encoding bits 0 and 1, one would require a polarization switch that will switch the polarization of an electromagnetic wave from approximately horizontal (encoding the bit 0) to approximately vertical (encoding the bit 1), or vice versa. Absolute polarization purity is not a concern for such applications. As long as the two polarizations are distinguishable from each other (i.e., they are approximately orthogonal), it will suffice.
Microwave polarization switches are however rare since they are usually much larger than the microwave wavelength (1-10 cm) and hence not integrable on a chip, although there are some exceptions [3, 4]. Consequently, most polarization switches work at optical frequencies where they are still larger than the optical wavelength (1 m) but small enough to be integrated on a chip. Optical polarization switches are implemented with various techniques such as rotating waveplates [5], Babinet-Soleil compensators [6], Berek rotary compensators [7], fiber coil polarization controllers [8], Faraday rotators [9], degree of polarization generators [10], lithium niobate electro-optics [11], liquid crystals [12], digital micromirrors [13], graphene metasurfaces [14] and on-chip photonic circuits [15, 16, 17]. All of them are typically much larger than the wavelength, but that is hardly a problem in the domain of optics where the wavelength is around only 1 m. It becomes a serious problem in the microwave frequency region where the wavelength can be several cm. Therefore, sub-wavelength polarization switches are very attractive for the microwave regime. Here, we demonstrate a microwave polarization switch antenna (working at 1-6 GHz) that is orders of magnitude smaller than the wavelength. This allows integration on a chip. That can have multiple applications, such as in quantum computing [3, 18], remote sensing [19] and, most importantly, aggressively miniaturized ultra-compact polarization division multiplexers and digital data transmitters in the microwave range. Additionally, it turns out that it may have other important applications, such as in secure communication in embedded applications (e.g., medically implanted devices, tiny stealthy drones for defense and crime-fighting), stealthy antennas to operate in hostile environments and physically unclonable functions (PUF) for antenna authetication and trust.
@@ 2 Results and Discussion
The structure of the polarization switch is shown in Fig. 1(a). It consists of a periodic two-dimensional array of magnetostrictive nanomagnets (made of cobalt) deposited on a piezoelectric substrate (128∘ Y-cut LiNbO3). The presence of both a magnetostrictive and a piezoelectric component effectively forms a two-phase multiferroic. This is called an “artificial multiferroic magnonic crystal” (AMMC) because of its periodicity. It is the same system that was used in ref. [21] for beam steering. Here, it is used for a different purpose, namely polarization switching. All antenna properties such as the scattering parameter spectrum, radiation efficiencies at different frequencies, etc. can be found in refs. [20, 21] and some are also repeated in the Supporting Information.
@@
The 128∘ Y-cut LiNbO3 substrate is 0.5 mm thick. The cobalt nanomagnets are slightly elliptical with major axis 110 nm and minor axis 100 nm. The thickness is 6 nm. Each nanomagnet has a 5 nm thick Ti layer underneath for adhesion to the substrate. The nanomagnet array covers an area of 100 m 100 m and is fabricated with electron-beam-lithography using a Raith Voyager e-beam writer. The substrate is spin-coated with a single layer of PMMA resist spun at 2500 rpm and baked at 110∘C for 2 minutes. It is then patterned with e-beam, and developed in a methyl isobutyl ketone and isopropyl alcohol (MIBK-IPA, 1:3) solution for 60 seconds, followed by a cold IPA rinse. After the patterning is complete, a 5 nm-thick titanium (Ti) adhesion layer is deposited using electron beam evaporation at a base pressure of 2.3 10-7 Torr, followed by the deposition of 6 nm thick cobalt. Lift-off is performed using remover PG solution. A scanning electron micrograph of the nanomagnets can be found in ref. [20]. The electrodes for launching surface acoustic waves are square patches delineated with optical lithography. They are made of Al which is 1 m thick and have lateral dimensions of 1.5 mm 1.5 mm.
@@
A surface acoustic wave (SAW) can be launched in two mutually orthogonal directions by activating electrode pairs (3,4) or (5,6) (see Fig. 1). The electrode pairs are square patches instead of interdigitated transducers (IDTs) since IDTs are also narrow band filters that do not allow launching of multiple SAW frequencies, which is required. In our case, the SAWs are not the usual Rayleigh waves, but a mixture of Rayleigh, Sezawa, Lamb modes. This does not matter since the only purpose of the SAW is to produce time varying strain in the magnetostrictive nanomagnets and make their magnetizations precess (via the Villari effect) to produce spin waves. The spin waves are confined (quantized) standing waves in the nanomagnets whose power and phase profiles were reported in [20, 22]. They radiate electromagnetic waves in space via magnon-photon coupling [20, 21, 23]. Here magnon-photon coupling implies that the spin waves transfer their energy to electromagnetic waves via mode coupling and this results in the emission of electromagnetic waves in space (antenna functionality). Unlike a conventional electromagnetic antenna that radiates electromagnetic waves owing to fluctuating charges or time-varying electric dipoles, these structures (nano-antennas) radiate electromagnetic waves due to fluctuating magnetization associated with spin waves [20, 21, 23, 24, 25, 26]. Hence, they do not behave as traditional antennas and are not constrained by the Harrington limit [27, 28] or Chu’s limit [29, 30] that afflict the gain and bandwidth of traditional antennas.
@@
In the past, we measured the radiation pattern of these samples by propagating a SAW in two different directions in the piezoelectric substrate [20]. We can apply a microwave frequency voltage between either electrode pairs (3,4) or (5,6) to launch SAWs in two mutually perpendicular directions. We then measure the radiation patterns in an anechoic chamber for these two different directions of SAW propagation. The patterns are measured in the plane of the nanomagnets and in the two planes transverse to the plane of the nanomagnets. They are measured in an AMS-8701 Anechoic Chamber, Antenna Measurement System using a 3164-10 Open Boundary Quad-ridged Horn Antenna. The sample (antenna) is always placed at a distance of 284.5 cm from the horn antenna which ensures that we are measuring the far-field radiation pattern at all frequencies.
@@
These radiation patterns can be found in ref. [20] and also the Supporting Information. They show significant anisotropy (radiation intensity is direction-dependent) at all frequencies measured (despite the antenna being an ultra-sub-wavelength structure that should behave like a point source which radiates isotropically), but more importantly, the radiation pattern depends on the direction of SAW propagation (the difference between “orientation 1” and “orientation 2”). The difference is, of course, more pronounced at some frequencies than at others. This latter feature is very likely caused by the fact that the strengths of magnon-phonon and magnon-photon couplings are frequency-dependent. A rigorous theory that can explain the frequency dependence is outside the scope of this work and will be explored in the future.
@@
We carried out micromagnetic simulations with OOMMF and MuMax3 to shed light on the directional dependence of the radiated intensity. These simulations (based on Landau-Lifshitz-Gilbert equations) yield the time evolution of the magnetization in the nanomagnets in the presence of periodic strain due to the SAW, and hence reveal the spin waves that are excited by the SAW. We found that the spin wave patterns are strongly influenced by whether the surface acoustic wave propagates parallel to the major axis of the elliptical nanomagnets (as would be the case if electrodes 3 and 4 in Fig. 1 were excited) or parallel to the minor axis of the nanomagnets (as would be the case if electrodes 5 and 6 were excited instead; see supplementary material of [21]). Since it is the spin waves that radiate the electromagnetic waves via magnon-photon coupling, the electromagnetic radiation pattern also depends strongly on the direction of SAW propagation [20].
@@
Why the spin wave pattern depends on the direction of SAW propagation is easily understood. The surface acoustic wave subjects the nanomagnets to periodic strain. Strain acts like an effective magnetic field in a magnetostrictive nanomagnet and in the case of SAW, an effective periodic magnetic field will appear in the nanomagnet [22]. This periodic field will make the magnetization process and oscillate in time, thereby producing a spin wave. The direction of the effective magnetic field is along the direction of SAW propagation [22]. Therefore, if we change the direction of SAW propagation, we will change the direction of the effective time-varying magnetic field in the nanomagnet and this will change the axis of precession of the magnetization within the nanomagnet, resulting in a change in the oscillations of the x-, y- and z-components of the magnetization, i.e. the spin wave pattern will change [21]. This then will change the emission intensity of electromagnetic waves in different directions, resulting in a change in the radiation pattern. Ref. [20] reported the dependence of the radiation pattern on the direction of SAW propagation for both horizontal and vertical polarizations, while ref.[21] found that changing the direction of SAW propagation changes the radiation spectrum as well. The latter is due to the fact that when the spin wave patterns change, their frequency components change too and that changes the spectrum of the emitted electromagnetic radiation.
@@
Changing the spin wave patterns in space by changing the direction of SAW propagation also changes the polarization of the radiated beam in any direction. This was not studied earlier and here we report that study. A phenomenolgical theory to explain why the polarization will depend on the direction of SAW propagation is provided in the Supporting Information.
@@ 2.1 Polarization dependence on the direction of SAW propagation
In Fig. 2, we first plot the radiation patterns in the plane of the nanomagnets for both vertical and horizontal polarizations in absolute units (as opposed to in dbi) for both directions of SAW propagation. This is done for different frequencies.
Let us say that at a given frequency the horizontal polarization value in a given direction is and the vertical polarization value is . We define a polarization angle as
@@
In Fig. 3, we show the radial plot of at different frequencies in radians. Note that at 5.9 GHz, in the 70∘ direction, is nearly radians for orientation 2 and nearly 0 radians for orientation 1 (see also Fig. s6 in the Supporting Information). Thus, by switching the direction of SAW propagation from orientation 1 to orientation 2, we can change the polarization of the beam radiated in the 70∘ direction at 5.9 GHz (in the nanomagnets’ plane) from nearly horizontal to nearly vertical. We can place the receiver at 70∘ angle as shown in the right panel of Fig. 3 and receive either horizontal polarization or vertical polarization by switching the direction of SAW propagation, i.e., by switching the electrode pairs that are activated to launch the SAW. This is the basis of the polarization switch for digital communication. We can transmit bits 0 and 1, encoded in horizontal and vertical polarizations, respectively, by switching between two pairs of electrodes. Note that the positions of the transmitting device and the receiver do not have to be fixed, but their relative alignment have to be fixed in this scheme. This enables point-to-point communication via polarization.
Take another frequency of 5.25 GHz and the 275∘ direction. For orientation 1, is very close to /2 (again see Fig. S6 in the Supporting Information), meaning that the emitted beam is vertically polarized whereas for orientation 2, is close to 0, meaning that the emitted beam is horizontally polarized. This is the opposite of the previous case because here orientation 1 corresponds to bit 1 and orientation 2 corresponds to bit 0. Thus, there are multiple directions and frequencies at which point to point communication is enabled via polarization encoding. This feature also lends itself to mutiple-input-multiple-output (MIMO) antenna functionality. Multiple bit streams, each with its own carrier frequency, can be transmitted in multiple different directions simultaneously to improve reliability, countering signal fading, etc.
The specific directions and the specific frequencies for communication (70∘ and 5.9 GHz or 275∘ and 5.25 GHz in this case) will depend on the antenna parameters such as size, shape, spacing, material composition, etc. of the nanomagnets and can be changed by changing these parameters. However, to develop a theory and simulation framework for designing the nanomagnet array to obtain a specific direction at a specific frequency is a major undertaking in itself and beyond the scope of this work. Fortunately, this is not needed for the applications mentioned here. One will simply measure the radiation pattern at different frequencies and experimentally obtain the correct directions and frequencies for communication. This is a calibration exercise – every antenna is calibrated. Once an antenna is calibrated, it is ready for use.
2.2 Covert communication
@@
The fact that the point-to-point communication via polarization encoding works only for a specific alignment of the transmitter and receiver makes covert communication [31, 32] possible. The authorized sender has access to the polarization radial plot and can inform the authorized receiver where to place the receiving device. An unauthorized receiver, intent on intercepting the message (eavesdropping), will not know where to place the receiver and will place it at a wrong location with very high probability. The eavesdropper will therefore not receive binary polarization states and hence cannot eavesdrop or intercept the message. This is similar to steganography [33, 34] and eliminates the need for encryption or cryptography which are always vulnerable to sophisticated attacks.
2.3 Physical unclonable function
@@
The scheme is also unclonable since the specific alignments and frequencies that work will vary slightly from sample to sample because of unavoidable manufacturing variations involved in the fabrication of the AMMC. Thus, the correct alignment direction and frequency are a “biometric” of any chosen antenna and hence can make the antenna act as a physical unclonable function (PUF) that cannot be reproduced or predicted [35], thereby providing strong security.
2.4 Stealthy antenna
Finally, the above feature introduces an element of stealth in the antenna operation. Correct bit streams are transmitted only the right direction at the right frequency and an eavesdropper whose receiver is not placed in the right location and tuned to the right frequency will receive garbled bit streams. For example, at 5.9 GHz, = 0 for both directions of SAW propagation (both orientations 1 and 2) in the 340∘ direction. Hence, a receiver placed in that 340∘ direction will receive a constant stream of 0’s if the horizontal polarization encodes the bit 0. This bit stream contains no message. Similarly, in the 165∘ direction, = radians for orientation 2 and 0.61 radians for orientation 1 at 5.9 GHz. Thus, the beam is 100% vertically polarized for the choice of orientation 2 while being 67% horizontally polarized and 33% vertically polarized for the choice of orientation 1 if the chosen frequency is again 5.9 GHz. The latter impure polarization states do not encode binary bits unambiguously and hence carry no meaningful message. Thus, the message is concealed from anyone without precise knowledge of the correct orientation and frequency. This enables stealth.